OpenStudy (cherrilyn):

use the integral test to determine if the infinite series is convergent or divergent

OpenStudy (cherrilyn):

\[\sum_{n=5}^{\infty} 1/\sqrt{n-4}\]

OpenStudy (anonymous):

\[\int\limits_{5}^{\infty} (x-4)^{-1/2} dx\]

OpenStudy (cherrilyn):

so it's convergent?